OpenStudy (anonymous):

If f(x)=(x+2/2) then what is f^-1(x)

OpenStudy (anonymous):

x^2/x+1

OpenStudy (anonymous):

\[f(x)=\frac{ x+2 }{ 2 } what is f ^{-1}(x)\]

OpenStudy (anonymous):

\[A. \frac{ 1 }{ 2 }x+1\]

OpenStudy (anonymous):

\[B. 2x+1\]

OpenStudy (anonymous):

\[C. 2x-2\]

OpenStudy (anonymous):

\[D. 2x+2\]

OpenStudy (anonymous):

put y = (x + 2)/2 now try to get x in terms of y , replace it with x and that is f ^-1 x

OpenStudy (anonymous):

Thanks! :)

OpenStudy (anonymous):

yw :)